OpenStudy (anonymous):

Medal and fan Carbonic acid and the corresponding bicarbonate has a pKa of 6.38. If the pH of the buffer was 6.05, which species would be in higher quantity? It would be the bicarbonate, correct?

OpenStudy (anonymous):

@aaronq

OpenStudy (anonymous):

@Abmon98

OpenStudy (aaronq):

use the Henderson-Hasslebalch equation until you are comfortable \(pH=pKa+log\dfrac{[A^-]}{[HA]}\) \(6.05=6.38+log\dfrac{[A^-]}{[HA]}\) \(\dfrac{[A^-]}{[HA]}=10^{(6.05-6.38)}=\dfrac{0.4677}{1}\)

OpenStudy (aaronq):

so there is \(\dfrac{0.4677}{1.4677}*100\%=31.9\%\) bicarbonate and 69.1 % carbonic acid.

OpenStudy (anonymous):

Why do you divide by 1.467 instead of 1?

OpenStudy (aaronq):

because \([A^-]=0.4677\) and \([HA]=1\) so to get the percent you have to divide by the total

OpenStudy (anonymous):

Oh okay thankyou!!

OpenStudy (aaronq):

no problem!